# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v2

@@ VIII.2 End-to-end simulations
We find cluster squeezing thresholds of $12.1,12.7,13$ dB for $20$, $15$ and $10$ PhANTM steps. We see that the required cluster squeezing increases as the number of PhANTM steps, and therefore the mean cat state amplitude $\alpha$, decreases. The lower panels in Fig. 13 show the mean and standard deviation of the GKP effective squeezing distributions output from the breeding process. We also compare the adaptive breeding protocol, used in Fig. 13, to standard breeding, where cat states must be squeezed exactly to the GKP spacing. With standard breeding, no threshold was found below $13.5$ dB, indicating that adaptive breeding can be used to reduce the threshold by at least $1.5$ dB and possibly more. While our simulation takes into account all generated GKP states, not just those above threshold, we can quantify the proportion of GKP states that fall above threshold, allowing us to not only compare the Gaussian squeezing resources required but also the generation probability with other GKP generation protocols. To this end, we determine the likelihood that a GKP state drawn randomly from the generated distribution has a squeezing level above the threshold, that is the combination of squeezing in each quadrature puts it in the ’correctable’ region of Fig. 12. This likelihood can then be taken as the success probability of our protocol. Some examples are shown in table 1.
@@
While direct simulations of the required Gaussian squeezing thresholds for other GKP generation methods are missing from the literature, we can compare to several approaches that benchmark themselves by generating $10$ dB GKP states, which are at the threshold for several fault-tolerant architectures Tzitrin et al. (2021); Fukui (2023); Fukui et al. (2018). The highest success probability for single-shot GKP generation protocols are found in Ref. Takase et al. (2024); however they require input $18-20$ dB of Gaussian squeezing, and photon number resolution up to 40 photons. On the other hand, lower input squeezing resources can be used as in Ref. Tzitrin et al. (2020), requiring $12$ dB, and Ref. Winnel et al. (2024) is as low as $6\mathrm{dB}$. However, in the former, success probabilities are on the order of $10^{-5}$ and the latter requires the generation of large, $n=10$, Fock states which are experimentally challenging .
IX Discussion
@@
The experimental hardware needed for the architecture is ambitious but feasible. The squeezing level needed is below the state-of-the-art levels achieved in free-space Vahlbruch et al. (2016). In recent years, the squeezing measured from on-chip sources has steadily improved, reaching $8.3$ dB Kashiwazaki et al. (2023), with efforts underway to reach levels seen in free-space. The other major resource needed in our approach is photon number resolving detection. While TES devices offer resolutions in excess of $100$ photons Eaton et al. (2022b) and have traditionally been used in CVQC, SMSPDs Kong et al. (2024) with low data acquisition requirements operating at higher temperatures have been shown to detect up to $10$ photons, sufficient for our architecture. Simultaneously, chip-integrated room-temperature PNR technologies capable of detecting up to $10$ photons at MHz scales are also in development Nehra et al. (2019). Finally, while losses in photonic circuit components need to be minimized for implementing our architecture on photonic chips, Silicon nitride platforms currently provide some of the lowest losses available for passive components Team (2024) at levels that look promising for FTQC.
@@
Multiple improvements, optimizations, and validations of the presented architecture remain to be explored. Higher probability methods for cat state generation via optimizing the squeezing and photon number resolution are possible. Adaptive breeding, used in both sensor state and magic state generation protocols, can be optimized with advanced adaptation algorithms. In the case of magic states, targeting a higher success probability while simultaneously reducing the quality of GKP states needed for magic generation looks possible with further development of the CV-QEC process with non-Gaussian states. The squeezing requirements of the architecture could also be lowered by moving to cluster states with a lower valency Paesani and Brown (2023) and employing more sophisticated GKP decoders, such as the ones taking into account correlations introduced by linear optics Walshe et al. (2024). While the effect of detection losses in PhANTM has been investigated Eaton et al. (2022a), estimating the loss and noise thresholds to target for hardware development would require more accurate and comprehensive error models, including the effects of Gaussian noise on effective squeezing and optical loss on Wigner negativity at each step of our architecture.
@@
While many improvements could be made and the performance under losses are yet to be fully quantified, we believe that the modular and passive architecture for photonic CV FTQC presented in the paper will constitute a foundational design for photonic quantum computers above the FT threshold in the future.
X Acknowledgements
Part of the QEC simulation code was funded under DARPA Contract HR00112390051. The views, opinions, and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.
@@ X Acknowledgements
We acknowledge insightful conversations with QC82 co-founders Olivier Pfister, Xu Yi, Joe Campbell, and Andreas Beling on theoretical and experimental CVQC, squeezing sources, homodyne detection, photon number resolution, and many other topics. We are thankful to Leandre Brunel for sharing some of his work on GKP lattice construction methods.

### v3

@@ 8.2 End-to-end simulations
We find thresholds for cluster squeezing needed for PhANTM and teleportation-based squeezing gate of $12.1$, $12.7$, and $13$ dB for $20$, $15$ and $10$ PhANTM steps, respectively. We see that the required cluster squeezing increases as the number of PhANTM steps, and therefore the mean cat state amplitude $\alpha$, decreases. The lower panels in Fig. 13 show the mean and standard deviation of the GKP effective squeezing distributions output from the breeding process. We also compare the adaptive breeding protocol, used in Fig. 13, to standard breeding, where cat states must be squeezed exactly to the GKP spacing. With standard breeding, no threshold was found below $13.5$ dB, indicating that adaptive breeding can be used to reduce the threshold by at least $1.5$ dB and possibly more. While our simulation takes into account all generated GKP states, not just those above threshold, we can quantify the proportion of GKP states that fall above threshold, allowing us to not only compare the Gaussian squeezing resources required but also the generation probability with other GKP generation protocols. To this end, we determine the likelihood that a GKP state drawn randomly from the generated distribution has a squeezing level above the threshold, that is the combination of squeezing in each quadrature puts it in the ’correctable’ region of Fig. 12. This likelihood can then be taken as the success probability of our protocol. Some examples are shown in table 1.
@@
While direct simulations of the required Gaussian squeezing thresholds for other GKP generation methods are missing from the literature, we can compare to several approaches that benchmark themselves by generating $10$ dB GKP states, which are at the threshold for several fault-tolerant architectures [27, 89, 84]. The highest success probability for single-shot GKP generation protocols are found in Ref. [57]; however they require input $18-20$ dB of Gaussian squeezing, and photon number resolution up to 40 photons. On the other hand, lower input squeezing resources can be used as in Ref. [49], requiring $12$ dB, and Ref. [63] is as low as $6\mathrm{dB}$. However, in the former, success probabilities are on the order of $10^{-5}$ and the latter requires the generation of large, $n=10$, Fock states which are experimentally challenging .
9 Discussion
@@
The experimental hardware needed for the architecture is ambitious but feasible. The squeezing level needed is below the state-of-the-art levels achieved in free-space [90]. In recent years, the squeezing measured from on-chip sources has steadily improved, reaching $8.3$ dB [91], with efforts underway to reach levels seen in free-space. The other major resource needed in our approach is photon number resolving detection. While TES devices offer resolutions in excess of $100$ photons [92] and have traditionally been used in CVQC, SMSPDs [30] with low data acquisition requirements operating at higher temperatures have been shown to detect up to $10$ photons, sufficient for our architecture. Simultaneously, chip-integrated room-temperature PNR technologies capable of detecting up to $10$ photons at MHz scales are also in development [93]. Finally, while losses in photonic circuit components need to be minimized for implementing our architecture on photonic chips, Silicon nitride platforms currently provide some of the lowest losses available for passive components [29] at levels that look promising for FTQC.
@@
Multiple improvements, optimizations, and validations of the presented architecture remain to be explored. Higher probability methods for cat state generation via optimizing the squeezing and photon number resolution are possible. Adaptive breeding, used in both sensor state and magic state generation protocols, can be optimized with advanced adaptation algorithms. In the case of magic states, targeting a higher success probability while simultaneously reducing the quality of GKP states needed for magic generation looks possible with further development of the CV-QEC process with non-Gaussian states. The squeezing requirements of the architecture could also be lowered by moving to cluster states with a lower valency [94] and employing more sophisticated GKP decoders, such as the ones taking into account correlations introduced by linear optics [82].
While the effect of losses in PhANTM has been investigated earlier [55], appendix 11.9 provides simulations and a discussion of the effect of losses on the adaptive breeding and QEC protocols. We briefly note that PhANTM is more sensitive to propagation and homodyne losses than to photon number resolution losses, and shows significant reduction in the Wigner negativity of cat states for losses beyond 1%. As for breeding, the effective squeezing of GKP states from breeding is reduced by less than $1$ dB with 3% homodyne loss. Similarly, the GKP effective squeezing needed to be at $10.2$ dB threshold for fault tolerance is increased to about $11.6$ dB with a 5% loss. This demonstrates that losses are less important for adaptive breeding and QEC than for cat generation through PhANTM. Estimating the loss and noise targets especially with further design improvements would require more comprehensive error models and simulations, including the effects of Gaussian noise on effective squeezing and optical loss on Wigner negativity at each step of the architecture. An end-to-end simulation incorporating losses is left for future work.
@@
While many improvements could be made and the performance under losses are yet to be fully quantified, we believe that the modular and passive architecture for photonic CV FTQC presented in the paper will constitute a foundational design for fault tolerant photonic quantum computers in the future.
10 Acknowledgements
@@
Part of the QEC simulation code was funded under DARPA Contract HR00112390051. The views, opinions, and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government. We acknowledge insightful conversations with QC82 co-founders Olivier Pfister, Xu Yi, Joe Campbell, and Andreas Beling on theoretical and experimental CVQC, squeezing sources, homodyne detection, photon number resolution, and many other topics. We are thankful to Leandre Brunel for sharing some of his work on GKP lattice construction methods.